# PiLLow: Enhancing Efficient Instruction Fine-tuning via Prompt Matching

### v1

@@ 1 Introduction
Therefore, can we attain a similar performance level to SFT by merely employing a comparable amount of resources as used by LoRA? Can we realize it using LoRA-fine-tuned LLMs’ reserved in-context learning (ICL) capacity? Our approach, named PiLLow, trains a Prompt matching net using ReInforcement Learning to improve fine-tuning LLMs under LOW-resource settings. We train an RL agent to select exemplars from a comprehensive “prompt set” which can be defined by users or split from the training data, and subsequently merge these examples to form a prompt which is then added at the beginning of the input text that is fed into LoRA model. Figure 1 shows an example where a 2-shot prompt is given by PiLLow agent based on the user input. Our approach becomes particularly beneficial in commercial applications where user input styles vastly deviate from those found in the pre-training corpus but are confined to limited variations. In these circumstances, the RL agent can efficiently learn to choose examples that best resonate with the specific query posed to the LLMs and therefore achieve comparable performance with direct SFT training. Our contributions can be summarized as follows:
@@ 2.1.2 In-Context Learning
In-context learning (ICL) is a method that enhances LLMs by supplying specific contexts using a handful of examples, or prompts, to steer the model’s behavior and produce the required results Dong et al. (2022). Efficient prompts can direct the model’s responses by offering pertinent information via sentences, keywords, instructions, or examples. ICL enables users to tailor the model for specialized tasks or fields, leveraging a relatively smaller dataset composed of examples and intended outputs. Nevertheless, as Zhao et al. (2021) emphasizes, ICL can be highly sensitive to the setup of prompts, encompassing prompt templates, in-context examples, and the order of the examples. (For more related work in prompting, please refer to Section A)
@@ 3.5 Results
Our observations in M-Score indicate that PiLLow significantly enhances the Excellent rate while reducing the Fail rate when compared to the LoRA model. This signifies a considerable improvement in quality in comparison to the LoRA model. Note that the 560m model trained with both PiLLow and LoRA does not improve that much and even degrades a little bit compared with the original pre-trained model, and on Alpaca PiLLow even performs worse than LoRA on the reward metric. However, there is no such problem for the 1b1 and 7b1 models. Therefore, we can conclude that our proposed PiLLow is particularly suitable for large-scale LLMs which inherently possess enhanced ICL and emergent capabilities. Importantly, the application of LoRA does not diminish these intrinsic abilities of the LLMs. We refer readers to Appendix D for example LLM inputs and outputs.
3.6 Ablation Study
@@ 3.6.1 Why RL?
While it may appear that the task introduced in this paper could be solved by simply matching the prompt most similar to the question during the inference phase, our first ablation study reveals that PiLLow outperform simple matching and LoRA model. This is due to the potential misalignment between the sentence encoder and LLM training data, which means the most semantically matched prompt might not yield the best answer. Additionally, as users may need to switch LLMs for different tasks which may require different objectives for LLMs outputs, PiLLow can optimize the outputs based on reward signals on multiple perspectives. We refer the readers to Appendix B for the experiment results and more details.
@@ 3.6.2 Number of Shots
We also investigate the impact of increasing the number of “shots”. Results indicate a slight improvement when the number of shots increases. However, using more shots may introduce irrelevant prompts that can disrupt the output of the LLMs. Additionally, a higher number of shots can reduce the PiLLow’s training efficiency. Hence, we recommend one prompt in practice. We refer the readers to Appendix C for the experiment results and more details.
@@ 4 Conclusion
Appendix A Automatic Prompting
Since writing prompts manually is time-consuming and cost-intensive, a number of methods have been proposed to automate the prompting process. In continuous prompting (a.k.a soft prompting) Liu et al. (2023a), prompting is performed directly in the embedding space of the language models. However, by their continuous nature, such prompts are not human-understandable. Prefix Tuning Li and Liang (2021) adds a sequence of continuous task-specific prompt embeddings to the front of input texts in each layer of LM while keeping the LM’s parameters frozen. Similarly, Prompt Tuning Lester et al. (2021) prepends the input texts with special tokens to form a template and directly tune the token embeddings without updating LM’s parameters. Unlike the two methods, P-Tuning Liu et al. (2021b) removes the restriction on adding the prompt embedding to the beginning of the input. They define that the prompt tokens can be inserted anywhere in the input sequence and can only be inserted in the input rather than any other model layer.
Appendix B Why RL?
RL techniques have been widely used in multiple industrial application and achieved significant improvement in numerous domain Qu et al. (2023); Coronato et al. (2020); Qiu et al. (2022); Chen et al. (2022); Hambly et al. (2023). Some may wonder why RL is even necessary in our settings since it seems that the tasks can be solved by simply matching the prompt with the largest similarity with the question. Here we conduct an ablation study on whether we use RL to solve the tasks. As can be seen in Table 2, with all else being equal, an RL-trained prompt matching network performs better than simple matching (SimMatch) which performs almost the same with pure LoRA. We attribute such a performance gap to the LLM in that the sentence encoder and the LLM may not be trained on the same data, so the prompt chosen by the encoder that best matches the question, i.e. seemed “matched” to humans, may not seem that “matched” to the LLM. Even if the chosen prompt is truly the most semantically similar one to the question, it may not best prompt to assist the LLM in generating an appropriate answer. In addition, users might want to switch to different LLMs to adapt to different downstream tasks, where such "appropriateness" might be defined differently, therefore an RL-based training framework is necessary.
Method C-Score PPL R/w M-Score
LoRA 3.09 78.81 5.83 0.09/0.21/0.70
SimMatch 3.12 79.26 5.81 0.12/0.26/0.62
PiLLow 3.21 67.36 5.89 0.14/0.39/0.47
Appendix C Number of Prompts
We ablate on the number of shots to study how the number of exemplars affects the performance. As can be seen in Table 3, as we increase the number of shots from 1 to 3, the ChatGPT score and reward increase a little, but in general, the measurements do not change too much. Intuitively, the more exemplars are given, the better the LLM output would be. However, it imposes difficulty to the matching net to ensure the selected exemplars are all helpful for prompting the LLM. It is possible that an irrelevant prompt is newly chosen and corrupts the LLM output. Also, a large number of shots makes the RL training slow. Therefore, we recommend using a small number of number of shots to balance the negative effects. In practice, we recommend just implement one prompt for PiLLow with best efficiency.
Type C-Score PPL R/w M-Score
1-shot 3.21 67.36 5.89 0.14/0.39/0.47
2-shot 3.19 66.95 6.05 0.14/0.41/0.45
3-shot 3.23 69.92 6.17 0.16/0.38/0.46
Appendix D Example LLM Input & Output

### v2

@@ 1 Introduction
Therefore, can we attain a similar performance level to SFT by merely employing a comparable amount of resources as used by LoRA? Can we realize it using LoRA-fine-tuned LLMs’ reserved in-context learning (ICL) capacity? Our approach, named PiLLow, trains a Prompt matching net using ReInforcement Learning to improve fine-tuning LLMs with LOW computation resources. We train an RL agent to select exemplars from a comprehensive “prompt set” which can be defined by users or split from the training data, and subsequently merge these examples to form a prompt which is then added at the beginning of the input text that is fed into LoRA model. Figure 1 shows an example where a 2-shot prompt is given by PiLLow agent based on the user input. Our approach becomes particularly beneficial in commercial applications where user input styles vastly deviate from those found in the pre-training corpus but are confined to limited variations. In these circumstances, the RL agent can efficiently learn to choose examples that best resonate with the specific query posed to the LLMs and therefore achieve comparable performance with direct SFT training. Our contributions can be summarized as follows:
@@ 2.1.2 In-Context Learning
In-context learning (ICL) is a method that enhances LLMs by supplying specific contexts using a handful of examples, or prompts, to steer the model’s behavior and produce the required results Dong et al. (2022). Efficient prompts can direct the model’s responses by offering pertinent information via sentences, keywords, instructions, or examples. ICL enables users to tailor the model for specialized tasks or fields, leveraging a relatively smaller dataset composed of examples and intended outputs. Nevertheless, as Zhao et al. (2021) emphasizes, ICL can be highly sensitive to the setup of prompts, encompassing prompt templates, in-context examples, and the order of the examples. (For more related work in prompting, please refer to Section 4)
@@ 3.5 Results
Our observations in M-Score indicate that PiLLow significantly enhances the Excellent rate while reducing the Fail rate when compared to the LoRA model. This signifies a considerable improvement in quality in comparison to the LoRA model. Note that the 560m model trained with both PiLLow and LoRA does not improve that much and even degrades a little bit compared with the original pre-trained model, and on Alpaca PiLLow even performs worse than LoRA on the reward metric. However, there is no such problem for the 1b1 and 7b1 models. Therefore, we can conclude that our proposed PiLLow is particularly suitable for large-scale LLMs which inherently possess enhanced ICL and emergent capabilities. Importantly, the application of LoRA does not diminish these intrinsic abilities of the LLMs. We refer readers to Appendix A for example LLM inputs and outputs.
3.6 Ablation Studies
@@ 3.6.1 Why RL?
RL techniques have been widely used in multiple industrial application and achieved significant improvement in numerous domain Qu et al. (2023); Coronato et al. (2020); Qiu et al. (2022); Chen et al. (2022); Hambly et al. (2023). Some may wonder why RL is even necessary in our settings since it seems that the tasks can be solved by simply matching the prompt with the largest similarity with the question. In this section, we conduct ablation study to show the superiority of RL in our setting.
As can be seen in Table 2, results reveal that PiLLow outperforms simple matching (SimMatch) and LoRA model. This is due to the potential misalignment between the sentence encoder and LLM training data—the most semantically matched prompt might not yield the best answer. The sentence encoder and the LLM may not be trained on the same data distribution, so the prompt chosen by the encoder that best matches the question, i.e. seemed “matched” to humans, may not seem that “matched” to the LLM. Even if the chosen prompt is truly the most semantically similar one to the question, it may not best prompt to assist the LLM in generating an appropriate answer. In addition, users might want to switch to different LLMs to adapt to different downstream tasks, where such “appropriateness" might be defined differently, therefore an RL-based training framework is necessary.
Method C-Score PPL R/w M-Score
LoRA 3.09 78.81 5.83 0.09/0.21/0.70
SimMatch 3.12 79.26 5.81 0.12/0.26/0.62
PiLLow 3.21 67.36 5.89 0.14/0.39/0.47
@@ 3.6.2 Number of Shots
We also investigate the impact of increasing the number of “shots”. We ablate on the number of shots to study how the number of exemplars affects the performance. As can be seen in Table 3, as we increase the number of shots from 1 to 3, the ChatGPT score and reward increase a little, but in general, the measurements do not change too much. Intuitively, the more exemplars are given, the better the LLM output would be. However, it imposes difficulty to the matching net to ensure the selected exemplars are all helpful for prompting the LLM. It is possible that an irrelevant prompt is newly chosen and corrupts the LLM output. Also, a large number of shots makes the RL training slow. Therefore, we recommend using a small number of number of shots to balance the negative effects. In practice, we recommend just implement one prompt for PiLLow with best efficiency.
Type C-Score PPL R/w M-Score
1-shot 3.21 67.36 5.89 0.14/0.39/0.47
2-shot 3.19 66.95 6.05 0.14/0.41/0.45
3-shot 3.23 69.92 6.17 0.16/0.38/0.46
4 Related Work
Since manually writing prompts is time-consuming and labor-intensive, a number of methods have been proposed to automate the prompting process. In continuous prompting (a.k.a soft prompting) Liu et al. (2023a), prompting is performed directly in the embedding space of the language models. However, by their continuous nature, such prompts are not human-understandable. Prefix Tuning Li and Liang (2021) adds a sequence of continuous task-specific prompt embeddings to the front of input texts in each layer of LM while keeping the LM’s parameters frozen. Similarly, Prompt Tuning Lester et al. (2021) prepends the input texts with special tokens to form a template and directly tune the token embeddings without updating LM’s parameters. Unlike the two methods, P-Tuning Liu et al. (2021b) removes the restriction on adding the prompt embedding to the beginning of the input. They define that the prompt tokens can be inserted anywhere in the input sequence and can only be inserted in the input rather than any other model layer.
@@ 5 Conclusion
Appendix A Example LLM Input & Output